# Solving Linear Systems of Equations with the Quantum HHL Algorithm: A Tutorial on the Physical and Mathematical Foundations for Undergraduate Students

### v1

@@ V Computational Complexity of the Algorithm
Algorithm analysis is a field of Computer Science dedicated to studying computational efficiency, focusing on how runtime scales with input size. Although elements such as hardware characteristics and programming style can influence performance, theoretical analysis abstracts away these factors and evaluates how many basic operations are executed Papadimitriou (2003). For example, time complexity, defined by a function $C(n)$, is used to characterize the count of these elementary operations as a function of $n$ Levitin (2003).
In this setting, an algorithm’s efficiency varies with the nature of the input and can be analyzed under three scenarios: the best case, which bounds the minimal possible runtime; the worst case, which establishes an upper bound on the time required; and the average case, which offers an expected performance for random inputs Papadimitriou (2003). Any algorithm to find the maximum in a set of $n$ elements requires at least n1n-1 comparisons, establishing a lower bound for the complexity of this task and showing that the standard algorithm is optimal Cormen (2012).
The most important problems in Computer Science include the class of NP-complete problems; the effective utility of quantum computing for real-world challenges depends on its potential to tackle such problems Shor (1997); Ekert and Jozsa (1996). Deutsch, Simon, and Shor were pivotal in developing quantum algorithms that exhibit exponential speedups over classical computation Jozsa (1998); Nielsen and Chuang (2010), and their theoretical breakthroughs attracted not only the scientific community but also significant industrial investment in quantum technologies. Given the importance of these aspects, in this section we discuss computational concepts related to the complexity of the HHL algorithm, highlighting its strengths compared to classical algorithms, as well as limitations and potential in practical applications. For a more detailed discussion of computational complexity, see Appendix B.
@@ V.1 Comparison with classical algorithms
The HHL algorithm emerged with the promise of an exponential advantage over classical algorithms for solving SLEs Harrow et al. (2009). This has drawn the attention of both the scientific and industrial communities, since the significant growth in data involved in SLEs underscores the importance of analyzing this algorithm to find solutions efficiently and in reduced time Alase et al. (2022). In some cases, a quantum computer can approximate a function of the solution in time logarithmic in $N$ and polynomial in the condition number $k=|\lambda_{\max}/\lambda_{\min}|$ of the matrix $A$ Aaronson (2015); Zaman et al. (2023); Duan et al. (2020a). Therefore, HHL can yield significant advantages in many configurations when $N$ is large and $k$ is small.
@@
For classical algorithms, the major obstacle in solving an SLE is matrix inversion, which also depends on $N$ and $k$. As $k$ grows, $A$ approaches a non-invertible matrix and solutions become less stable Diblík et al. (2016). The HHL algorithm assumes that the singular values of $A$ satisfy $k^{-2}I\leq A^{\dagger}A\leq I$ Harrow et al. (2009), which would imply a runtime scaling as $k^{2}\log(N)/\epsilon$, where $\epsilon$ is the additive error in the output state $|x\rangle$ Zaman et al. (2023). In this sense, HHL’s greatest advantage over classical methods occurs when $k$ and $1/\epsilon$ are polynomial in $\log(N)$, achieving an exponential speedup Harrow et al. (2009).
@@
For comparison, the conjugate gradient method—one of the triumphs of classical computing—can solve an SLE with linear complexity in $N$ under suitable conditions Dervovic et al. (2018). If $A$ is sufficiently sparse, with $s$ nonzero entries per row, and positive definite, then this method uses $O(\sqrt{k}\log(1/\epsilon))$ matrix–vector multiplications, each costing $O(sN)$, totaling $O(Ns\sqrt{k}\log(1/\epsilon))$ operations. If $A$ is negative definite, $O(k\log(1/\epsilon))$ multiplications are needed, yielding $O(Nsk\log(1/\epsilon))$ total Harrow et al. (2009). In more pessimistic scenarios, classical methods can reach $O(N^{3})$ using Gaussian elimination Zudio et al. (2016) and $O(N^{2.33})$ using block Krylov techniques via recursive low-displacement-rank factorizations Peng and Vempala (2021).
@@
On the other hand, HHL shows that a quantum circuit built with $n$ qubits and $T$ gates can be simulated by inverting an $O(1)$-sparse matrix $A$ of dimension $N=O(2^{n}k)$ Harrow et al. (2009). The condition number $k$ is $O(T^{2})$ when $A$ is positive and $O(T)$ otherwise Harrow et al. (2009). This implies that a classical algorithm running in time $poly(\log N,k,1/\epsilon)$ could simulate a quantum algorithm with $poly(n)$ gates in time $poly(n)$ Zaman et al. (2023). Taking these complexities into account, HHL achieves a number of operations on the order of $O(\log(N)\,s^{2}\,k^{2}/\epsilon)$ Harrow et al. (2009).
@@
In general, the classical computational complexity for solving an SLE with a dense $N\times N$ matrix is $O(N^{3})$, as in Gaussian elimination, although block-Krylov-based factorizations can reduce this to $O(N^{2.33})$ Peng and Vempala (2021). However, when the matrix is sparse and has favorable properties such as symmetry and positive definiteness, classical algorithms like conjugate gradient can approach $O(N)$ complexity under ideal conditions Dervovic et al. (2018); Shewchuk (1994).
@@ V.2 Algorithm Limitations
Recent research has underscored the importance of careful analysis when it comes to practical applications of quantum algorithms Aaronson (2015); Tang (2022); McClean et al. (2018); Cerezo et al. (2024); Gil-Fuster et al. (2024). In this sense, understanding critical aspects of the algorithm involves analyzing not only its strengths but also its inherently theoretical limitations. In the case of HHL, part of these limitations were discussed earlier in terms of the conditions required for the algorithm to work Harrow et al. (2009). In this subsection, we explore other limitations that can pose challenges for applying the algorithm in real scenarios.
@@
A first issue—common to many quantum algorithms—is the preparation of the state $\left|b\right\rangle$ as in Eq. (6). In classical simulations, state preparation is a straightforward routine: one applies a set of rotation gates to obtain a normalized vector (see Box 3). However, the number of gates required grows exponentially with the vector size, making preparation infeasible for larger instances and significantly increasing the computational cost of the algorithm Sambhaje and Chaurasia (2024). In practice, it is necessary to use a Quantum Random Access Memory (QRAM), which uses qubits to access any quantum superposition stored in its cells Giovannetti et al. (2008a).
@@
Several works have investigated architectures capable of enabling efficient applications Giovannetti et al. (2008b); Zidan et al. (2021); Xu et al. (2023), while others have criticized their feasibility Jaques and Rattew (2023); Steiger (2016). Discussions surrounding QRAM question whether a genuinely quantum architecture can be built, the implementation cost in terms of energy and computational resources, and qubit decoherence times Jaques and Rattew (2023). All these constraints make preparing an arbitrary state in HHL difficult. Thus, one alternative to this problem is to prepare a state with approximately uniform amplitudes, which significantly reduces the steps required for storing states in a QRAM or may even allow quantum algorithms to be used without directly relying on it [^2] Generating an equal (or nearly equal) superposition can obviate the need for QRAM to load arbitrary amplitudes. This choice simplifies state preparation, since it requires a number of gates that grows polynomially, instead of configuring each coefficient individually Giovannetti et al. (2008b).. Note that this already introduces another limitation for the algorithm—the state $\left|b\right\rangle$ would cease to be arbitrary. Nevertheless, even under this condition, a classical algorithm can perform inner-product operations with complexity $\log n/\epsilon$ when the vector is uniformly distributed Aaronson (2015). In that case, HHL would no longer be genuinely useful.
@@
Henceforth, suppose QRAM is available. The next step in the algorithm is to apply QPE, where another difficulty arises precisely in the stage of applying the unitary operators $e^{iAt}$ for different values of $t$. If the matrix $A$ is sufficiently sparse, then applying $e^{iAt}$ scales linearly for $s\ll n$ Berry et al. (2015). In the best cases, $A$ is sparse enough to achieve an exponential advantage, as demonstrated in Clader et al. (2013); Lloyd et al. (2016).
@@
Now suppose the algorithm has been executed successfully, without the obstacles previously mentioned; even so, there remains a final issue: measuring the state $\left|x\right\rangle$. As illustrated in Fig. 5, the quantity of interest in the algorithm lies in the probability amplitudes of $\left|x\right\rangle$. This information can be retrieved through repeated measurements of the relevant qubit(s), in a process known as state tomography, which allows reconstruction of its probability distribution. In Box 10, we present only measurement in the computational basis, since the goal is to analyze the relationship between the states $\left|01\right\rangle$ and $\left|11\right\rangle$. The $1024$ shots indicate that the circuit is executed $1024$ times Nielsen and Chuang (2010); Mermin (2007), but a complete reconstruction of the state $\left|x\right\rangle$ requires additional measurements in other bases. In general, the number of steps required for this process (standard tomography with Pauli-basis measurements) is on the order of $N^{c}$, where $N$ is the number of qubits and $c$ is a problem-dependent constant, such as circuit depth, measurement scheme, and so on. This fact again limits HHL’s exponential advantage Aaronson (2015).
@@ V.3 Algorithm Potential
The previous section discussed the main limitations of the HHL algorithm, such as the need for tomography to explicitly extract the components of the solution vector and the difficulty of dealing with arbitrary initial states. However, HHL also presents strengths that justify its central role in other quantum algorithms. Indeed, HHL’s potential does not lie in providing an explicit solution to a linear system, but in preparing a quantum state corresponding to the solution, which can be efficiently leveraged in different contexts Zaman et al. (2023).
One of the main applications of HHL is computing expectation values of observables with respect to the quantum solution vector. When the objective is not to know the components of $\left|x\right\rangle$ directly, it is still possible to obtain $\langle x|M|x\rangle$ for some observable $M$, thereby providing a substantial quantum advantage over classical approaches Zheng et al. (2025). This ability to efficiently compute global properties has applications across diverse problem classes, with emphasis on Hamiltonian simulations Zheng et al. (2025); Baskaran et al. (2023).
Moreover, HHL naturally integrates as a subroutine within more complex quantum algorithms Sambhaje and Chaurasia (2024); Zheng et al. (2025); Liu et al. (2022); Jing et al. (2024); Baskaran et al. (2023). In quantum machine learning—such as quantum neural networks or quantum data classification—HHL is used to solve SLEs that arise in intermediate steps Duan et al. (2020b); Liu et al. (2022). In such cases, the quantum state $\left|x\right\rangle$ resulting from HHL is directly used as input for subsequent stages, without the need to measure all components explicitly, preserving the quantum advantage of the procedure. In other scenarios, HHL’s initial state $\left|b\right\rangle$ need not be prepared from scratch, as it may be the output of other quantum routines Clader et al. (2013). This feature simplifies preparation and reduces the associated cost, enabling practical use of the algorithm across multiple stages of state manipulation and computation Liu et al. (2022).
@@ Appendix A Introduction to Quantum Computing
Quantum computing is a new computational paradigm that seeks to apply concepts from quantum mechanics to computing by exploiting purely quantum properties such as superposition and entanglement Nielsen and Chuang (2010); Jesus et al. (2021); Galvão et al. (2024). In particular cases, it is possible to establish analogous relationships between the two models of computation, classical and quantum Mermin (2007). In classical computing, the fundamental element is the bit (BInary digiT), a unit of information that takes binary values $\{0,1\}$ Mano (1993). In quantum computing, by contrast, bits carry quantum properties and are therefore called qubits (QUantum BITs). Qubits are two-level systems that, unlike classical bits, can exist in linear superpositions of their basis states. One of the most common representations used in the study of quantum computing to describe these systems is the use of bra and ket vectors, introduced by Dirac’s notation Dirac (1981).
Dirac notation provides a compact way to represent inner products between two states $\alpha$ and $\beta$ with $\left\langle\alpha\middle|\beta\right\rangle$. While bras are row vectors, written as $\left\langle\alpha\right|$, kets are written as $\left|\beta\right\rangle$. In the computational basis, qubits initially assume the states $\left|0\right\rangle$ or $\left|1\right\rangle$, analogous to classical bits, and these states form the orthonormal basis of the quantum space Nielsen and Chuang (2010). Eq. (61) shows the vector representation corresponding to these basic qubit states.
@@
Analogous to classical computing, quantum computing requires logic gates, characterized by unitary operators that guarantee conservation of the system’s probability Nielsen and Chuang (2010). These operators can be described by matrices which, when applied to the qubits defined above, change their states. In general, such matrices have size $2^{n}\times 2^{n}$ for $n$ qubits on which the operation will be applied. The most common cases are $n=1$ square matrices, such as the Hadamard gate
@@
These elements constitute the structural basis of a quantum circuit, usually described within the formalism of digital quantum computation Mermin (2007), as illustrated in Fig. 7. The dynamics of a quantum circuit are characterized by a sequence of unitary operators (quantum logic gates) applied to individual qubits or entangled states, followed by projective measurements that extract classical information. This architecture enables the implementation of quantum algorithms through coherent manipulation and accurate readout of the qubits’ final states.

### v2

@@ V Computational Complexity of the Algorithm
Algorithm analysis is a field of Computer Science dedicated to studying computational efficiency, focusing on how runtime scales with input size. Although elements such as hardware characteristics and programming style can influence performance, theoretical analysis abstracts away these factors and evaluates how many basic operations are executed papadimitriou2003computational. For example, time complexity, defined by a function $C(n)$, is used to characterize the count of these elementary operations as a function of $n$ levitin2003.
In this setting, an algorithm’s efficiency varies with the nature of the input and can be analyzed under three scenarios: the best case, which bounds the minimal possible runtime; the worst case, which establishes an upper bound on the time required; and the average case, which offers an expected performance for random inputs papadimitriou2003computational. Any algorithm to find the maximum in a set of $n$ elements requires at least n1n-1 comparisons, establishing a lower bound for the complexity of this task and showing that the standard algorithm is optimal cormen2012algoritmos.
The most important problems in Computer Science include the class of NP-complete problems; the effective utility of quantum computing for real-world challenges depends on its potential to tackle such problems Shor_1997; ekert1996quantum. Deutsch, Simon, and Shor were pivotal in developing quantum algorithms that exhibit exponential speedups over classical computation Jozsa_1998; nielsen2010quantum, and their theoretical breakthroughs attracted not only the scientific community but also significant industrial investment in quantum technologies. Given the importance of these aspects, in this section we discuss computational concepts related to the complexity of the HHL algorithm, highlighting its strengths compared to classical algorithms, as well as limitations and potential in practical applications. For a more detailed discussion of computational complexity, see Appendix B.
@@ V.1 Comparison with classical algorithms
The HHL algorithm emerged with the promise of an exponential advantage over classical algorithms for solving SLEs harrow2009quantum. This has drawn the attention of both the scientific and industrial communities, since the significant growth in data involved in SLEs underscores the importance of analyzing this algorithm to find solutions efficiently and in reduced time PhysRevResearch.4.023237. In some cases, a quantum computer can approximate a function of the solution in time logarithmic in $N$ and polynomial in the condition number $k=|\lambda_{\max}/\lambda_{\min}|$ of the matrix $A$ aaronson2015read; zaman; duan2020survey. Therefore, HHL can yield significant advantages in many configurations when $N$ is large and $k$ is small.
@@
For classical algorithms, the major obstacle in solving an SLE is matrix inversion, which also depends on $N$ and $k$. As $k$ grows, $A$ approaches a non-invertible matrix and solutions become less stable Diblík2016Exponential. The HHL algorithm assumes that the singular values of $A$ satisfy $k^{-2}I\leq A^{\dagger}A\leq I$ harrow2009quantum, which would imply a runtime scaling as $k^{2}\log(N)/\epsilon$, where $\epsilon$ is the additive error in the output state $|x\rangle$ zaman. In this sense, HHL’s greatest advantage over classical methods occurs when $k$ and $1/\epsilon$ are polynomial in $\log(N)$, achieving an exponential speedup harrow2009quantum.
@@
For comparison, the conjugate gradient method—one of the triumphs of classical computing—can solve an SLE with linear complexity in $N$ under suitable conditions dervovic2018quantum. If $A$ is sufficiently sparse, with $s$ nonzero entries per row, and positive definite, then this method uses $O(\sqrt{k}\log(1/\epsilon))$ matrix–vector multiplications, each costing $O(sN)$, totaling $O(Ns\sqrt{k}\log(1/\epsilon))$ operations. If $A$ is negative definite, $O(k\log(1/\epsilon))$ multiplications are needed, yielding $O(Nsk\log(1/\epsilon))$ total harrow2009quantum. In more pessimistic scenarios, classical methods can reach $O(N^{3})$ using Gaussian elimination sscad and $O(N^{2.33})$ using block Krylov techniques via recursive low-displacement-rank factorizations peng2021solvingsparselinearsystems.
@@
On the other hand, HHL shows that a quantum circuit built with $n$ qubits and $T$ gates can be simulated by inverting an $O(1)$-sparse matrix $A$ of dimension $N=O(2^{n}k)$ harrow2009quantum. The condition number $k$ is $O(T^{2})$ when $A$ is positive and $O(T)$ otherwise harrow2009quantum. This implies that a classical algorithm running in time $poly(\log N,k,1/\epsilon)$ could simulate a quantum algorithm with $poly(n)$ gates in time $poly(n)$ zaman. Taking these complexities into account, HHL achieves a number of operations on the order of $O(\log(N)\,s^{2}\,k^{2}/\epsilon)$ harrow2009quantum.
@@
In general, the classical computational complexity for solving an SLE with a dense $N\times N$ matrix is $O(N^{3})$, as in Gaussian elimination, although block-Krylov-based factorizations can reduce this to $O(N^{2.33})$ peng2021solvingsparselinearsystems. However, when the matrix is sparse and has favorable properties such as symmetry and positive definiteness, classical algorithms like conjugate gradient can approach $O(N)$ complexity under ideal conditions dervovic2018quantum; 10.5555/865018.
@@ V.2 Algorithm Limitations
Recent research has underscored the importance of careful analysis when it comes to practical applications of quantum algorithms aaronson2015read; tang2022dequantizing; McClean2018BarrenPI; cerezo2023does; gil2024relation. In this sense, understanding critical aspects of the algorithm involves analyzing not only its strengths but also its inherently theoretical limitations. In the case of HHL, part of these limitations were discussed earlier in terms of the conditions required for the algorithm to work harrow2009quantum. In this subsection, we explore other limitations that can pose challenges for applying the algorithm in real scenarios.
@@
A first issue—common to many quantum algorithms—is the preparation of the state $\left|b\right\rangle$ as in Eq. (6). In classical simulations, state preparation is a straightforward routine: one applies a set of rotation gates to obtain a normalized vector (see Box 3). However, the number of gates required grows exponentially with the vector size, making preparation infeasible for larger instances and significantly increasing the computational cost of the algorithm sambhaje2024hhl. In practice, it is necessary to use a Quantum Random Access Memory (QRAM), which uses qubits to access any quantum superposition stored in its cells PhysRevLett.100.160501.
@@
Several works have investigated architectures capable of enabling efficient applications giovannetti2008architectures; zidan2021novel; xu2023systems, while others have criticized their feasibility jaques2023qram; pirsa_PIRSA:16080019. Discussions surrounding QRAM question whether a genuinely quantum architecture can be built, the implementation cost in terms of energy and computational resources, and qubit decoherence times jaques2023qram. All these constraints make preparing an arbitrary state in HHL difficult. Thus, one alternative to this problem is to prepare a state with approximately uniform amplitudes, which significantly reduces the steps required for storing states in a QRAM or may even allow quantum algorithms to be used without directly relying on it [^2] Generating an equal (or nearly equal) superposition can obviate the need for QRAM to load arbitrary amplitudes. This choice simplifies state preparation, since it requires a number of gates that grows polynomially, instead of configuring each coefficient individually giovannetti2008architectures.. Note that this already introduces another limitation for the algorithm—the state $\left|b\right\rangle$ would cease to be arbitrary. Nevertheless, even under this condition, a classical algorithm can perform inner-product operations with complexity $\log n/\epsilon$ when the vector is uniformly distributed aaronson2015read. In that case, HHL would no longer be genuinely useful.
@@
Henceforth, suppose QRAM is available. The next step in the algorithm is to apply QPE, where another difficulty arises precisely in the stage of applying the unitary operators $e^{iAt}$ for different values of $t$. If the matrix $A$ is sufficiently sparse, then applying $e^{iAt}$ scales linearly for $s\ll n$ berry2015simulating. In the best cases, $A$ is sparse enough to achieve an exponential advantage, as demonstrated in PhysRevLett.110.250504; lloyd2014quantum.
@@
Now suppose the algorithm has been executed successfully, without the obstacles previously mentioned; even so, there remains a final issue: measuring the state $\left|x\right\rangle$. As illustrated in Fig. 5, the quantity of interest in the algorithm lies in the probability amplitudes of $\left|x\right\rangle$. This information can be retrieved through repeated measurements of the relevant qubit(s), in a process known as state tomography, which allows reconstruction of its probability distribution. In Box 10, we present only measurement in the computational basis, since the goal is to analyze the relationship between the states $\left|01\right\rangle$ and $\left|11\right\rangle$. The $1024$ shots indicate that the circuit is executed $1024$ times nielsen2010quantum; mermin2007quantum, but a complete reconstruction of the state $\left|x\right\rangle$ requires additional measurements in other bases. In general, the number of steps required for this process (standard tomography with Pauli-basis measurements) is on the order of $N^{c}$, where $N$ is the number of qubits and $c$ is a problem-dependent constant, such as circuit depth, measurement scheme, and so on. This fact again limits HHL’s exponential advantage aaronson2015read.
@@ V.3 Algorithm Potential
The previous section discussed the main limitations of the HHL algorithm, such as the need for tomography to explicitly extract the components of the solution vector and the difficulty of dealing with arbitrary initial states. However, HHL also presents strengths that justify its central role in other quantum algorithms. Indeed, HHL’s potential does not lie in providing an explicit solution to a linear system, but in preparing a quantum state corresponding to the solution, which can be efficiently leveraged in different contexts zaman.
One of the main applications of HHL is computing expectation values of observables with respect to the quantum solution vector. When the objective is not to know the components of $\left|x\right\rangle$ directly, it is still possible to obtain $\langle x|M|x\rangle$ for some observable $M$, thereby providing a substantial quantum advantage over classical approaches zheng2024early. This ability to efficiently compute global properties has applications across diverse problem classes, with emphasis on Hamiltonian simulations zheng2024early; baskaran2022adapting.
Moreover, HHL naturally integrates as a subroutine within more complex quantum algorithms sambhaje2024hhl; zheng2024early; Liu_2022; JING2024122878; baskaran2022adapting. In quantum machine learning—such as quantum neural networks or quantum data classification—HHL is used to solve SLEs that arise in intermediate steps DUAN2020126595; Liu_2022. In such cases, the quantum state $\left|x\right\rangle$ resulting from HHL is directly used as input for subsequent stages, without the need to measure all components explicitly, preserving the quantum advantage of the procedure. In other scenarios, HHL’s initial state $\left|b\right\rangle$ need not be prepared from scratch, as it may be the output of other quantum routines PhysRevLett.110.250504. This feature simplifies preparation and reduces the associated cost, enabling practical use of the algorithm across multiple stages of state manipulation and computation Liu_2022.
@@ Appendix A Introduction to Quantum Computing
Quantum computing is a new computational paradigm that seeks to apply concepts from quantum mechanics to computing by exploiting purely quantum properties such as superposition and entanglement nielsen2010quantum; jesus2021computaccao; galvao2024possibilidades. In particular cases, it is possible to establish analogous relationships between the two models of computation, classical and quantum mermin2007quantum. In classical computing, the fundamental element is the bit (BInary digiT), a unit of information that takes binary values $\{0,1\}$ mano1993computer. In quantum computing, by contrast, bits carry quantum properties and are therefore called qubits (QUantum BITs). Qubits are two-level systems that, unlike classical bits, can exist in linear superpositions of their basis states. One of the most common representations used in the study of quantum computing to describe these systems is the use of bra and ket vectors, introduced by Dirac’s notation dirac1981principles.
Dirac notation provides a compact way to represent inner products between two states $\alpha$ and $\beta$ with $\left\langle\alpha\middle|\beta\right\rangle$. While bras are row vectors, written as $\left\langle\alpha\right|$, kets are written as $\left|\beta\right\rangle$. In the computational basis, qubits initially assume the states $\left|0\right\rangle$ or $\left|1\right\rangle$, analogous to classical bits, and these states form the orthonormal basis of the quantum space nielsen2010quantum. Eq. (61) shows the vector representation corresponding to these basic qubit states.
@@
Analogous to classical computing, quantum computing requires logic gates, characterized by unitary operators that guarantee conservation of the system’s probability nielsen2010quantum. These operators can be described by matrices which, when applied to the qubits defined above, change their states. In general, such matrices have size $2^{n}\times 2^{n}$ for $n$ qubits on which the operation will be applied. The most common cases are $n=1$ square matrices, such as the Hadamard gate
@@
These elements constitute the structural basis of a quantum circuit, usually described within the formalism of digital quantum computation mermin2007quantum, as illustrated in Fig. 7. The dynamics of a quantum circuit are characterized by a sequence of unitary operators (quantum logic gates) applied to individual qubits or entangled states, followed by projective measurements that extract classical information. This architecture enables the implementation of quantum algorithms through coherent manipulation and accurate readout of the qubits’ final states.